# OTAD: An Optimal Transport-Induced Robust Model for Agnostic Adversarial Attack

### v1

@@ V-C Model comparison and settings
We empirically compare OTAD with three types of defense methods, i.e., adversarial training, adversarial purification, and Lipschitz networks. For the adversarial training methods, we choose the commonly used PGD [30] and FGSM [10] adversarial training. We evaluate the methods by unseen threats and thus use $l_{\infty}$ adversarial training except for industrial data. The noise level of the $l_{\infty}$ adversarial training $\epsilon$ is set to $\epsilon=0.1$ on MNIST, $\epsilon=8/255$ on CIFAR10, $\epsilon=4/255$ on TinyImageNet. For the adversarial purification methods, we choose median filter [34], STL [35], GAN [40] and DiffPure [42] to represent image pre-processing, optimization-based reconstruction, generative model-based purification models, respectively. When using the BPDA + PGD attack on these purification methods, we approximate the purification process using the identity map during back-propagation. For the non-purification methods, BPDA + PGD attack reduces to standard PGD. For the Lipschitz networks, we compare with $l_{\infty}$-distance net [24, 25] and a 1-Lipschitz $l_{2}$ network SOC+, which incorporates skew orthogonal convolutions with certificate regularization and Householder activation [44]. The hyperparameters of these models are the same as the original papers.
@@ V-D1 Performance on MNIST
FGSM adversarial training 72.7 0 0 0
@@
To further investigate the ability of OTAD, we vary the hyper-parameters $L-l$ and record the standard accuracy, robust accuracy, local Lipschitz constant of OTAD, and relative error (RE). The local Lipschitz constant is computed at a noise level $\epsilon$ around $0.3$ that equals the adaptive attack. RE measures the gap between ResNet $R(\cdot)$ and OTAD feature $\tilde{f}(\cdot)$,
@@
We can see that $L-l$ controls the local Lipschitz constant of OTAD (Fig. 2b). Thus, we may set smaller $L-l$ to make OTAD more robust. However, smaller $L-l$ is not always better. The local Lipschitz constant of the original ResNet is approximately $5.6$. Smaller OTAD Lip ($<5.6$) results in better robustness (smaller disparity between standard and robust accuracy) but worse standard accuracy (Fig. 2a). That is because OTAD deviates from the ground truth. A larger OTAD Lip results in worse robustness, larger RE, and worse standard accuracy (Fig. 2c). That is because OTAD deviates from the ResNet. The best result can be achieved with $L-l$ slightly smaller than the empirical Lipschitz constant of the original ResNet.
Figure 2: Performance of OTAD with different hyperparameter $L-l$ against BPDA + PGD ($\epsilon=3$). (a) The increasing disparity between standard and robust accuracy as $L-l$ varies. (b) The local Lipschitz constant of OTAD (OTAD Lip) increases with $L-l$. (c) The relative error increases with increasing $L-l$.
@@
FGSM adversarial training 72.7 0 0 0
@@
FGSM adversarial training 72.7 0 0
@@ V-D2 Performance on the single-cell transcriptomics data
OTAD is suitable for other data types. We use the MERFISH retina single-cell transcriptomics data by Chen et al. [63] and preprocess it with Scanpy. The dataset is a 7-class cell-type classification problem containing 82122 single cells (65697 for training and 16425 for testing), and each cell contains expression profiles for 500 genes. Since it is hard to use generative models (adversarial purification) for single-cell transcriptomics data, we only compared OTAD with the PGD adversarial training (Table IV). OTAD still achieves better robust accuracy than the PGD adversarial training.
@@ V-E OTAD-T can deal with more complex data
For more complex datasets such as CIFAR10 and TinyImageNet, we test OTAD-T with DML-net for $K$ neighbor searching with parameters $K=10$, $L=2$, $l=0$, and $\delta^{1}=\delta^{2}=1$. As with OTAD, we first choose a gradient-free attack, BPDA + PGD, with a noise level of $\epsilon=0.5$ on both datasets. For CIFAR10, we focus on two types of DML-nets, i.e., a 3-block ResNet and an attention block. The results show that OTAD-T with a 3-block ResNet achieves the best accuracy, and OTAD-T with an attention block demonstrates robustness but has lower standard accuracy (Table VI).
@@
PGD adversarial training 66.8 52.3
Median filter 79.1 7.7
STL 78.3 14.9
APE-GAN 77.9 18.3
DiffPure 79.5 69.8
$l_{\infty}$-dist net 56.1 17.6
SOC+ 77.0 43.4
OTAD-T with attention 61.3 54.4
OTAD-T with ResNet 88.4 81.5
OTAD-T-NN with attention 62.2 54.1
OTAD-T-NN with ResNet 88.4 81.5
Introducing DML-net brings additional robustness challenges. We use an untargeted PGD attack to disrupt the learned features and test the performance of OTAD-T under this DML attack. Table VII shows the robustness of OTAD-T with different DML-nets under the DML attack on CIFAR10. Despite OTAD-T with ResNet’s strong defense against BPDA attacks, it can still be vulnerable to DML attacks. For a non-robust DML-net, we face a trade-off, i.e., better neighborhoods but increased vulnerability.
TABLE VII: Performance of different DML-nets in OTAD-T on CIFAR10 against DML attack ($\epsilon=0.5$).
@@
3-block ResNet 88.4 3.3
2-layer CNN 60.0 4.8
1-block ViT 53.1 32.8
An attention block 61.3 48.1
We also show the effectiveness of OTAD-T on TinyImageNet. The ResNet-18 is chosen for DML-net. Table VIII shows the results compared with other defense methods, OTAD-T with ResNet-18 achieves the best accuracy. Moreover, a precise DML-net can even help improve the standard accuracy of OTAD-T, outperforming the original network. The cooperation of the two networks achieves better results and it is difficult to attack when DML-net is unknown.
TABLE VIII: Performance of defense methods on TinyImageNet against BPDA + PGD ($\epsilon=0.5$).
@@
PGD adversarial training 43.0 26.5
Median filter 43.3 10.4
STL 41.0 12.8
APE-GAN 40.9 16.2
DiffPure 46.1 21.4
$l_{\infty}$-dist net 24.0 0.9
SOC+ 31.6 16.3
OTAD-T with ResNet 65.2 60.9
V-F OTAD-T-NN achieves fast inference
We design an end-to-end 6-block Transformer encoder (CIP-net) for fast inference (OTAD-T-NN). The number of tokens in CIP-net equals $12K+$ 1, where $K$ ($K=10$) is the number of neighbors. The training set of CIP-net is given by the solutions of the QCP solver of OTAD-T with attention. Thus, CIP-net is a neural network designed for this specific optimization problem. OTAD-T-NN achieves a much faster inference speed. We compare the average inference time of OTAD-T and OTAD-T-NN per sample on CIFAR10 (Table IX), and a 3-block ResNet is chosen for DML-net.
TABLE IX: Average inference time (standard deviation) per sample on CIFAR10 (seconds)
Method Inference time
OTAD-T 1.6436 ($4.36\times 10^{-3}$)
OTAD-T-NN 0.0047 ($1.59\times 10^{-8}$)
Besides efficient inference, Table VI shows the performance of OTAD-T-NN against BPDA + PGD remains robust. Additionally, though the CIP-net is trained on solutions from OTAD-T with attention, OTAD-T-NN with ResNet remains highly efficient.
OTAD-T-NN is now differentiable. We can use gradient-based adversarial attacks, such as PGD, to test the robustness of OTAD-T-NN. Table X shows the performance of defense methods against PGD attack. We do not test the STL and median filter due to the non-differentiable nature of their purification processes. Experiments show that OTAD-T-NN remains robust against PGD, indicating that gradient obfuscation is not the main reason for the robustness of OTAD.
TABLE X: Performance of defense methods on CIFAR10 against PGD ($\epsilon=0.5$).
Method Standard Acc Robust Acc
PGD adversarial training 66.8 52.3
APE-GAN 77.9 0.10
DiffPure 79.5 62.1
$l_{\infty}$-dist net 56.1 17.6
SOC+ 77.0 43.4
OTAD-T-NN with attention 62.2 60.2
OTAD-T-NN with ResNet 88.4 88.0
@@ V-F OTAD-T-NN achieves fast inference
We compared the robustness against the PGD attack of OTAD-T-NN trained by different values of $\alpha$ in Table XI. As $\alpha$ decreases, the robustness of the model also decreases, indicating that OTAD-T-NN’s robustness is primarily derived from the convex integration problem.
TABLE XI: Effect of $\alpha$ on the Robustness of OTAD-T-NN on CIFAR10 against PGD ($\epsilon=0.5$).
@@ V-G Compared to KNN
The KNN classifier predicts the class of a sample based on the majority class of its $K$ nearest training samples, which is robust. OTAD integrates information from neighboring points to compute a robust interpolation for the test data, functioning as a form of weighted KNN. Here, we compare OTAD and OTAD-T with KNN on MNIST and CIFAR10, respectively, while $L=2$ in OTAD and $L=12$ in OTAD-T. All methods search neighbors by the $l_{2}$ distance. For test data with $K$ neighbors, the KNN feature of the test point is derived from the mean of the features of these $K$ neighbors, while the OTAD feature is obtained by solving the convex integration problem. Table XII shows the standard accuracy of these methods with different $K$s. Although OTAD outperforms KNN by only an average of 0.83% on MNIST, OTAD-T outperforms KNN by an average of 7.7% on CIFAR10 across $K=$ 5, 10, 15. That demonstrates that OTAD shows distinctly higher performance.

### v2

@@ V-C Model comparison and settings
We empirically compare OTAD with three categories of defense methods, i.e., adversarial training, adversarial purification, and Lipschitz networks. OTAD’s two-step process derives a Lipschitz model from a standard ResNet or ViT, so we include existing Lipschitz networks for comparison. We also evaluate OTAD against classic adversarial defense methods to demonstrate its practical performance, including adversarial training methods that may exhibit reduced robustness on unseen threats and adversarial purification methods that modify the forward propagation like OTAD.
For adversarial training methods, we consider the primary PGD [30] adversarial training, FGSM [10] adversarial training, TRADES [32], MART [34], and AWP [35]. We evaluate these methods under two experimental protocols. First, we evaluate robustness under unseen threat models. Specifically, adversarial training methods (except for the industrial data) are trained with $l_{\infty}$ norm perturbations and evaluated against $l_{2}$ attacks. Since OTAD and other methods are not trained by adversarial samples, this setting corresponds to an unseen attack scenario for all methods. The noise level of the $l_{\infty}$ adversarial training $\epsilon$ is set to $\epsilon=0.1$ on MNIST and $\epsilon=8/255$ on CIFAR10. Second, we evaluate adversarial training methods under a norm-consistent setting, i.e., training and testing are conducted under the same perturbation norm (both $l_{2}$ norms, $\epsilon=0.5$ on CIFAR10). The training attacks run 40 iterations on MNIST and 20 iterations on CIFAR10. All methods are trained from the same base models in OTAD. Due to the hardness of adversarial training with ViT architecture, we also train ResNet-18 adversarially on CIFAR10 for comparison.
For adversarial purification methods, we choose median filter [36], STL [37], GAN [42], and DiffPure [44] to represent image pre-processing, optimization-based reconstruction, and generative model-based purification models, respectively. The classifier used in these purification methods is the same as in OTAD. When using the BPDA + PGD attack to these purification methods, we approximate the purification process with the identity map during back-propagation.
For Lipschitz networks, we compare with $l_{\infty}$-distance net [24, 25] and a 1-Lipschitz $l_{2}$ network SOC+, which incorporates skew orthogonal convolutions with certificate regularization and Householder activation [46]. The hyperparameters of these models are the same as the original papers. Specifically, the block size of SOC+ is 1 on MNIST and 3 on CIFAR10.
@@ V-D1 Performance on MNIST
FGSM adversarial training 97.7 0 0 0
@@
[…]
@@
We can see that LlL-l controls the local Lipschitz constant of OTAD (Fig. 2b). Thus, we may set smaller LlL-l to make OTAD more robust. However, smaller LlL-l is not always better. The local Lipschitz constant of the original ResNet is approximately $5.6$. Smaller OTAD Lip (<5.6<5.6) results in better robustness (smaller disparity between standard and robust accuracy) but worse standard accuracy (Fig. 2a). That is because OTAD deviates from the ground truth. A larger OTAD Lip results in worse robustness, larger RE, and worse standard accuracy (Fig. 2c). That is because OTAD deviates from the ResNet. The best result can be achieved with LlL-l slightly smaller than the empirical Lipschitz constant of the original ResNet.
Figure 2: Performance of OTAD with different hyperparameter LlL-l against BPDA + PGD ($\epsilon=3$). (a) The increasing disparity between standard and robust accuracy as LlL-l varies. (b) The local Lipschitz constant of OTAD (OTAD Lip) increases with LlL-l. (c) The relative error increases with increasing LlL-l.
@@
FGSM adversarial training 97.7 4.0 2.2 1.3
@@
FGSM adversarial training 97.7 0 0
@@ V-D2 Performance on the single-cell transcriptomics data
OTAD is suitable for other data types. We use the MERFISH retina single-cell transcriptomics data by Chen et al. [64] and preprocess it with Scanpy. The dataset is a 7-class cell-type classification problem containing 82122 single cells (65697 for training and 16425 for testing), and each cell contains expression profiles for 500 genes. Since it is hard to use generative models (adversarial purification) for single-cell transcriptomics data, we only compare OTAD with the PGD adversarial training (Table IV). OTAD still achieves better robust accuracy than the PGD adversarial training.
@@ V-E OTAD-T can deal with more complex data
For more complex datasets such as CIFAR10 and ImageNet, we test OTAD-T with DML-net for $K$ neighbor searching with parameters $K=10$, $L=2$, $l=0$, and $\delta^{1}=\delta^{2}=1$. As with OTAD, we first choose a gradient-free attack, BPDA + PGD. For CIFAR10, we focus on two types of DML-nets, i.e., a 3-block ResNet and an attention block. The results show that OTAD-T with a 3-block ResNet achieves the best accuracy, and OTAD-T with an attention block demonstrates robustness but has lower standard accuracy (Table VI). Additionally, OTAD-T maintains robustness against BPDA + PGD for multiple values of epsilon (Table VII).
When an effective DML-net is used, OTAD-T demonstrates strong robustness against BPDA+PGD. We evaluate the adversarial training methods with both ViT and ResNet-18, which achieve comparable robust accuracy across architectures. However, adversarial training methods are vulnerable to unseen threats, causing a significant drop in robustness compared to OTAD-T. In the norm-consistent setting, adversarial training methods achieve higher robust accuracy than in unseen threat evaluation. Nevertheless, their performance remains below that of OTAD-T, highlighting the superior robustness of our approach. Adversarial purification methods using powerful generative models, such as diffusion models, also exhibit strong robustness. OTAD-T achieves competitive results when compared to DiffPure. Importantly, OTAD does not require training an additional generative model and can be directly applied to different types of data. Lipschitz networks tend to perform modestly on more complex datasets, often resulting in lower standard accuracy. Therefore, while Lipschitz networks can be effective in certain tasks like MNIST, OTAD-T can deal with more complex data.
@@
Adversarial Training with $l_{\infty}$
PGD adversarial training 77.7 63.4
PGD adversarial training (ResNet-18) 82.6 62.4
TRADES 84.4 64.3
TRADES (ResNet-18) 88.6 64.3
MART 71.0 61.3
MART (ResNet-18) 85.2 64.9
TRADES + AWP 84.5 65.4
TRADES + AWP (ResNet-18) 89.7 67.1
Adversarial Training with $l_{2}$
PGD adversarial training 85.7 61.4
PGD adversarial training (ResNet-18) 89.7 66.3
TRADES 88.8 64.1
TRADES (ResNet-18) 90.2 70.2
MART 85.2 62.2
MART (ResNet-18) 89.5 71.6
TRADES + AWP 86.4 65.5
TRADES + AWP (ResNet-18) 89.4 74.1
Adversarial Purification
Median filter 89.5 34.2
STL 87.2 37.8
APE-GAN 84.3 30.1
DiffPure 88.1 83.5
Lipschitz Networks
$l_{\infty}$-dist net 56.1 20.3
SOC+ 77.8 45.9
Ours
OTAD-T with attention 60.6 59.0
OTAD-T with ResNet 91.2 86.1
OTAD-T-NN with attention 61.9 60.4
OTAD-T-NN with ResNet 91.1 85.8
TABLE VII: Performance of OTAD-T on CIFAR10 against BPDA+PGD with different perturbation magnitudes.
Method Standard Acc Robust Acc
$\epsilon=0.5$ $\epsilon=1.0$ $\epsilon=1.5$
OTAD-T with ResNet 91.2 86.1 82.9 79.9
OTAD-T with attention 60.6 59.0 58.2 57.9
Introducing DML-net brings additional robustness challenges. We use an untargeted PGD attack to disrupt the learned features and test the performance of OTAD-T under this DML attack. Table VIII shows the robustness of OTAD-T with different DML-nets under the DML attack on CIFAR10. Despite OTAD-T with ResNet’s strong defense against BPDA attacks, it can still be vulnerable to DML attacks. For a non-robust DML-net, we face a trade-off, i.e., better neighborhoods but increased vulnerability. Moreover, a precise DML-net can even help improve the standard accuracy of OTAD-T, outperforming the original network. The cooperation of the two networks achieves better results and it is difficult to attack when DML-net is unknown.
TABLE VIII: Performance of different DML-nets in OTAD-T on CIFAR10 against DML attack ($\epsilon=0.5$).
@@
3-block ResNet 91.2 9.5
2-layer CNN 58.3 9.8
1-block ViT 49.3 35.8
An attention block 60.6 49.6
OTAD-T requires neighborhood search in the training set to solve the CIP problem. Therefore, one of the main challenges for OTAD-T to scale to large datasets is the memory and computational cost associated with searching the training data. To address this, for large datasets like ImageNet, we may select or randomly sample a subset of data points in the inference stage to reduce memory and computational cost.
In each class of the training data, we randomly choose subsets of varying sizes, such as 50 (base), 10 (small), 5 (tiny), and 1 (nano), for the neighborhood search process in the inference stage. This results in 50K, 10K, 5K, and 1K samples for each corresponding subset size. We choose the ViT-B/16 pre-trained on ImageNet as DML-net. Based on the results in Table IX, the size of the training subset directly impacts OTAD-T’s performance, but the base version is effective enough. We can increase the subset size for better performance without being concerned about memory and computational cost. Thus, OTAD is effective on large dataset like ImageNet.
Additionally, the inference time does not significantly depend on the subset size at current scale. We decompose the OTAD process into three main components: neighborhood search, LP solver, and QCP solver. The neighborhood search step, which involves identifying K neighborhoods and feeding them into the CIP solver or CIP-net, is negligible in terms of computational cost at current scale. QCP solver is the most time-consuming step (Table X), with a substantial variance in execution time across different subset sizes. The neighborhood search step may affect the inference speed of OTAD, when larger subsets are used and more neighborhoods need to identify. Nevertheless, we can randomly select a training subset to ensure that OTAD can scale to large datasets, making it suitable for large-scale applications without significant performance degradation.
TABLE IX: Performance of OTAD-T with different sizes of subsets on ImageNet against BPDA + PGD ($\epsilon=2.0$).
OTAD-T Standard Acc Robust Acc
Base 78.8 68.1
Small 78.0 67.6
Tiny 68.4 56.2
Nano 24.2 21.8
TABLE X: Comparison of average inference time (standard deviation) per sample (seconds) and memory usage (GB) for OTAD-T with different size of subsets on ImageNet across the different computation steps.
Step Base Small Tiny Nano
Neighbor search 0.0219 s ($6.46\times 10^{-4}$) 0.0172 s ($2.43\times 10^{-4}$) 0.0172 s ($2.64\times 10^{-4}$) 0.0208 s ($5.70\times 10^{-4}$)
LP solver 1.6111 s (0.6714) 0.5325 s (0.0895) 0.6235 s (0.0965) 1.4162 s (0.4713)
QCP solver 9.5429 s (9.9035) 11.5611 s (26.7888) 9.1533 s (10.4927) 11.7192 s (25.4457)
Total 11.1765 s (10.7814) 12.1116 s (26.9571) 9.7947 s (10.6045) 13.1567 s (26.2487)
Memory Usage 62.2 GB 12.3 GB 8.6 GB 6.3 GB
V-F OTAD-T-NN achieves fast inference
We design an end-to-end 6-block Transformer encoder (CIP-net) for fast inference (OTAD-T-NN). The number of tokens in CIP-net equals $2K+1$, where $K$ ($K=10$) is the number of neighbors. Each token is mapped to a $d$-dimensional embedding ($d=2048$ on CIFAR10 and $d=4096$ on ImageNet). The training set of CIP-net is given by the solutions of the QCP solver of OTAD-T. Thus, CIP-net is a neural network designed for this specific optimization problem. OTAD-T-NN achieves a much faster inference speed. The additional computational cost is associated with training this Transformer, including making the training dataset and the training process. We compare the average inference time of OTAD-T and OTAD-T-NN per sample on CIFAR10 and ImageNet (Table XI). For CIFAR10, the 3-block ResNet is chosen for DML-net, while for ImageNet, the base version of OTAD-T is used.
TABLE XI: Average inference time (standard deviation) per sample on CIFAR10 and ImageNet (seconds).
Dataset Method Inference time
CIFAR10 OTAD-T 1.9640 s ($1.13\times 10^{-1}$)
OTAD-T-NN 0.0046 s ($2.66\times 10^{-3}$)
ImageNet OTAD-T 11.1765 s ($1.08\times 10^{1}$)
OTAD-T-NN 0.0138 s ($8.46\times 10^{-5}$)
Besides efficient inference, Table VI shows the performance of OTAD-T-NN against BPDA + PGD remains robust. Additionally, though the CIP-net is trained on solutions from OTAD-T with ResNet, OTAD-T-NN with attention remains highly efficient.
OTAD-T-NN is now differentiable. We can use gradient-based adversarial attacks, such as AutoAttack, to test the robustness of OTAD-T-NN. Table XII and XIII show the performance of defense methods against AutoAttack. We do not test the STL and median filter due to the non-differentiable nature of their purification processes. Similar to the results under BPDA + PGD, OTAD-T-NN remains robust and achieves best performance against one of the strongest adversarial attacks, AutoAttack, indicating that gradient obfuscation is not the main reason for the robustness of OTAD.
TABLE XII: Performance of defense methods on CIFAR10 against AutoAttack ($\epsilon=0.5$).
@@ V-F OTAD-T-NN achieves fast inference
Adversarial Training with $l_{\infty}$
PGD adversarial training 77.7 58.1
PGD adversarial training (ResNet-18) 82.6 57.9
TRADES 84.4 61.8
TRADES (ResNet-18) 88.6 59.3
MART 71.0 54.7
MART (ResNet-18) 85.2 60.1
TRADES + AWP 84.5 62.4
TRADES + AWP (ResNet-18) 89.7 62.5
Adversarial Training with $l_{2}$
PGD adversarial training 85.7 59.2
PGD adversarial training (ResNet-18) 89.7 63.0
TRADES 88.8 62.4
TRADES (ResNet-18) 90.2 67.0
MART 85.2 59.5
MART (ResNet-18) 89.5 68.8
TRADES + AWP 86.4 64.1
TRADES + AWP (ResNet-18) 89.4 71.7
Adversarial Purification
APE-GAN 84.3 0.0
DiffPure 88.1 74.3
Lipschitz Networks
$l_{\infty}$-dist net 56.1 1.0
SOC+ 77.8 41.8
Ours
OTAD-T-NN with attention 61.9 31.3
OTAD-T-NN with ResNet 91.1 76.3
TABLE XIII: Performance of OTAD-T-NN with different sizes of subsets on ImageNet against AutoAttack ($\epsilon=2.0$).
OTAD-T Standard Acc Robust Acc
Base 78.8 63.6
Small 77.4 58.6
Tiny 65.0 16.5
Nano 23.6 0.1
@@
We compare the robustness against AutoAttack of OTAD-T-NN trained by different values of $\alpha$ in Table XIV. As $\alpha$ decreases, the robustness of the model also decreases, indicating that OTAD-T-NN’s robustness is primarily derived from the convex integration problem.
TABLE XIV: Effect of $\alpha$ on the Robustness of OTAD-T-NN on CIFAR10 against AutoAttack ($\epsilon=0.5$).
@@ V-G Compared to KNN
The KNN classifier predicts the class of a sample based on the majority class of its $K$ nearest training samples, which is robust. OTAD integrates information from neighboring points to compute a robust interpolation for the test data, functioning as a form of weighted KNN. Here, we compare OTAD and OTAD-T with KNN on MNIST and CIFAR10, respectively, while $L=2$ in OTAD and $L=12$ in OTAD-T. All methods search neighbors by the $l_{2}$ distance. For test data with $K$ neighbors, the KNN feature of the test point is derived from the mean of the features of these $K$ neighbors, while the OTAD feature is obtained by solving the convex integration problem. Table XV shows the standard accuracy of these methods with different $K$s. Although OTAD outperforms KNN by only an average of 0.83% on MNIST, OTAD-T outperforms KNN by an average of 6.2% on CIFAR10 across $K=$ 5, 10, 15. That demonstrates that OTAD shows distinctly higher performance.